# A Tale of Two DL Cities: When Library Tests Meet Compiler

### v1

@@ VIII Conclusion
In this work, we propose Opera, a test-migration-based technique, to test the model loading stage in a lightweight manner (compared to grammar-based testing techniques with substantial implementation effort for supporting DL models under various DL libraries). Opera uses the human-written and tool-generated test inputs from DL libraries as migration sources. Then, Opera extracts the operator instances from DL library test inputs and wraps them to DL models as the test inputs for DL compilers. To improve the testing efficiency, Opera designs a diversity-based test prioritization strategy in order to detect more bugs within a given time budget. By applying Opera to eight frontends of three popular DL compilers (i.e., TVM, TensorRT, and OpenVINO), Opera detected $170$ previously unknown bugs (including $90$ confirmed bugs). The diversity-based test prioritization strategy in Opera achieves the average improvement of 11.9$\sim$47.4compared to general test prioritization in terms of APFD.
@@ Acknowledgments
We thank all the anonymous reviewers for their thoughtful and constructive comments on this work. This work was supported by the National Natural Science Foundation of China (Grant Nos. 62322208, 62232001), CCF Young Elite Scientists Sponsorship Program (by CAST), and Hong Kong Research Grant Council/General Research Fund (Grant No. 16205722).

### v2

@@ VIII Conclusion
In this work, we propose Opera, a migration-based technique, to test the model loading stage in a lightweight manner. Opera uses tests documented in DL libraries, and the tests generated by two recent fuzzers as migration sources. Then, Opera extracts the operator instances from DL library tests and wraps them based on templates into DL models as migrated tests. To improve the testing efficiency, Opera includes a diversity-based test prioritization strategy. By applying Opera to eight frontends of three popular DL compilers (i.e., TVM, TensorRT, and OpenVINO), Opera detected $170$ previously unknown bugs (including $90$ confirmed bugs). The diversity-based test prioritization strategy in Opera achieves the average improvement of 11.9$\sim$47.4compared to general test prioritization in terms of APFD.
@@ Acknowledgments
We thank all the anonymous reviewers for their thoughtful and constructive comments on this work. This work was supported by the National Natural Science Foundation of China (Grant Nos. 62322208, 12411530122, 62232001), CCF Young Elite Scientists Sponsorship Program (by CAST), and Hong Kong Research Grant Council/General Research Fund (Grant No. 16205722).